evening shift 08/21/99

John Shields (jshields@klik)
Sat, 21 Aug 1999 23:58:34 -0500

Shift Summary for eve shift, 08/21/99

On Shift: J. Shields, V. O'Dell

Hours of beam available at all intensities: 8.0
Hours of beam available at nominal intensity: 7.0
Hours of E832 data (special runs not included): 7.7
Hours of special runs (laser scans, muon runs...): 0.05

Accounting of lost time:

leaks in Tevatron magnet cooling system, switching polarization of
NM4AN, tape swapping

========================================================================
Runs ended this shift:
Run 14292 - 23.6M events - BEAM_832
Run 14293- BEAM_MUON
Run 14294 - 7.6M events - BEAM_832
Run 14295-97 - 0 events (Viv had to re-initialize a couple of times
before the correct magnet reading came up on
the screen during the run start-up process)

Tapes brought to FCC this shift:
(including tapes for online split)

QKC,QKS,QKK 297-299 & QKB104
QKC,QKS,QKK 300-302 & QKB105

on-line split: QKN026

========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:
E832 datataking, muon run,brief magnet cooling problems, minor shift
personnel snafu w.r.t magnet polarity for muon run, more E832 data-taking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14292 385 17:35 HCC inefficient-- stripe in HCC illumination
plot. Reloaded Etotal.

400 17:55 Tape warning: "BOY, I'M FULL!"

18:13 Finished swapping tapes, brought tapes QKC,QKS,QKK
300-302 & QKB105 to Feynman Center.

14293 1 18:30 begin muon run on new tapes

18:31 JS returns from FC, notices that tapes from PREV
shift still on control room table, and makes
another trip to FCC with QKC,QKS,QKK 297-299 & QKB104

18:32 Stopped muon run

14294 1 18:40 begin BEAM_832 run

19:15 DN puts 2 new bottle of Ar/Eth on line on east side

34 19:30 ktevana errors. Kfrended.

19:48 the ever-vigilant shift crew notices a red button
on Slow-DAQ. Ktsad.exe process has stopped on
atev2. Viv restarts it.

21:25 Viv realizes she flipped NM4AN the wrong direction
(ie from NEG to NEG) and panics! Luckily, fate
was on her side and beam chooses that moment to
die. Viv seizes the opportunity to flip the magnet.

21:27 cooling system for magnets has a leak. Beam down.
Estimate is 20 min.

21:45 Beam back, but magnet flipping still in progress.

14298 1 21:55 begin new run.

14298 42 22:49 JS swaps in a new CO2 bottle for TRDs.

23:35 chan 1104 has stab RMS of 4.86% Gone next
spill-- false alarm

89 23:53 ktevana errors. Kfrended.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift